# Machine Learning Should Maximize Welfare, Not (Only) Accuracy

### v1

@@ 3.3 Order 2: Enabling User Choices
Even if a system has the capacity to make decisions about users, often those very users will also have some say regarding final outcomes. For example, even in hiring and admissions, candidates much first choose to apply, and if selected, choose to accept. Thus, both reward for the system and utility to users depend on outcomes that result from the interplay of system decisions and user choices. Similarly to system decisions, we will model user choices as also depending on predictions, or more generally on the learned model $h$. Let ${c}_{{\pi}_{h}}(x)$ be the choice of user $x$ in response to $h$. Incorporating user choices into the system’s objective gives:
@@
  • Matching: Prime examples in which user choices drive economic outcomes are recommendation systems and online market platforms. These typically include a single learner whose role is to match users with relevant items (Jagadeesan et al., 2023b); once matched, it is the choices of users that ultimately determine allocations. Since matchings are often based on predictions, their quality determines outcomes for users, suppliers, and the system. This makes them unique in relation to conventional two-sided markets (Rochet and Tirole, 2003). In modern human-facing systems, key resources primarily include the attention, time, and budgets of users. This calls for careful and responsible modeling of user choice behavior (Kleinberg et al., 2024).
  • Competition: For many tasks and services of interest, users can now choose between multiple platforms or providers. This creates competition: once providers aim to maximize their market share, users themselves become the scarce resource, over which providers compete (Ben-Porat and Tennenholtz, 2019; Jagadeesan et al., 2023a; Yao et al., 2023, 2024). When the benefit of users depends on the quality of predictions, competition will revolve around which provider learns more accurate models—and for which users. This creates a novel ‘market for accuracy’, in which learning can have much impact on efficiency and equity.
@@
4 Alternative Views
While it is easy to agree that learning systems should be designed to enable the promotion of social good, there will likely be disagreement as to how. We propose to adopt the perspective of welfare economics, but there certainly exist alternative viewpoints and complementary approaches.
Give accuracy time.
One perspective is that if we give data enough time to accumulate and new methods enough time to improve, then machine learning will organically overcome the challenges we discussed. One example to draw on is how despite many advances in optimization, the simple gradient descent algorithm still drives most modern tools. Another is how large language models have demonstrated that simply predicting the next word with sufficient accuracy and on enough data gives rise to emergent phenomena far beyond this basic task. Our position is that limited resources is an inherent problem of any social system, whether technology-driven or not. We believe that scarcity should be addressed explicitly—but of course we may be proven wrong.
Divide and conquer.
Even if machine learning as a standalone solution does not suffice, one could argue that an economic approach can be applied on top of existing learning tools, rather than integrated within them as we propose. Hence, learning and policy can be advanced independently and combined only later. This is reasonable, and independent efforts and application will likely be required regardless of whether direct integration works or not. But there is increasing evidence that this will not suffice; in fact, the field of fair machine learning rose in response to the clear need for embedding social considerations within the learning objective itself. Advances in the study of fairness in learning have also shown that fairness constraints alone cannot guarantee equity, such as when learning effects accumulate over time (e.g., Liu et al., 2018) or as a result of strategic user behavior (Horowitz et al., 2024). We take these to suggest that the novelty in the interface between learning and economics requires a wholistic approach specialized for this intersection.
Welfare without welfare.
Welfare economics is not the only approach for reasoning about and facilitating welfare, nor is it free from issues and limitations in itself. Criticism includes its subjective nature; the need to measure and compare utility across individuals; the emphasis on cardinal rather than ordinal utilities; the reliance on assumptions of rational behavior; the susceptibility to externalities and other sources of market failure; the need for a centralized social planner entity; and challenges in policy evaluation. Other schools of thought in economics offer alternatives: For example, Sen’s capabilities approach (Sen, 1999) focuses on ensuring people are capable of achieving what they seek, rather than the value of what they obtain. Within machine learning, there been have calls for alternative approaches as well, such as to ‘democratize’ the issue of alignment using social choice theory (Conitzer et al., 2024; Ge et al., 2024; Fish et al., 2024). We view these as complementary to ours, and believe there is merit advancing welfare in machine learning simultaneously along several fronts.
@@ 5.3 Evaluation
As we state in Sec. 1.1, our focus has been on prediction policies (Kleinberg et al., 2015) that specify actions as $a={\pi}_{h}(x)={\pi}(h(x))$. The key assumption of prediction policies is that knowing the true $y$ suffices for making an optimal decision; hence the motivation for replacing $y$ with the prediction ${\hat{y}}=h(x)$. But in realistic settings, $y$ is rarely the only source of uncertainty. Most decision settings have a casual aspect, meaning that actions $a$ can affect outcomes $y$. Hence, there is no ‘true’ $y$ on which we can draw, and the optimal action $a$ (or ‘intervention’) must include an understanding of its causal effects on $y$.
@@ 6 Concluding Discussion
Economics studies the tensions that arise from having limited means but unlimited wants. Our main thesis is that while the appetite for what machine learning can give us only grows, we do not yet have the tools to integrate the limitations inherent to any social system in a way that supports informed policy choices. Welfare economics provides general guidelines for designing systems that enable a social planner to express and promote outcomes that align with societal preferences, as they relate to scarcity in resources. We have identified ways in which we believe this perspective applies to, and can be useful for, the design of learning systems that operate in social contexts and with human agency. Scarcity in learned systems is often far more illusive than in classic economic settings that consider tangible goods. But nonetheless, it exists, and its implications bear much significance for the value that we can hope to obtain from using machine learning as individuals. Our proposal to augment conventional learning frameworks with a notion of social welfare serves as a first step towards this goal. But the road ahead is long and challenging, and requires careful deliberation of the societal role we envision for machine learning.
Impact statement
This paper proposes that machine learning can and should be used to maximize social welfare. In principle, and by construction, the impact of our proposed framework on society aims to be positive. But our paper also points to the inherent difficulties of identifying, and making formal, what ‘good for society’ is. We lean on the field of welfare economics, which has for decades contended with this challenge, for ideas on how the learning community can begin to approach this daunting task. However, even if these ideas are conceptually appealing, the path to practical welfare improvement presents many challenges—some expected, others unforseen. For example, we may specify incorrect social welfare functions; or we may specify them correctly but be unable to optimize them appropriately; or we may be able to optimize but find that our assumptions are wrong, that theory differs from practice, or that there were other considerations and complexities that we did not take into account. For this we can look to other related fields—such as fairness, privacy, and alignment in machine learning—which have taken (and are still taking) similar journeys, and learn from both their success and mistakes.
Any discipline that seeks to affect policy should do so with much deliberation and care. Whereas welfare economics was designed with the explicit purpose of supporting (and influencing) policymakers, machine learning has found itself in a similar position, but likely without any planned intent. On the one hand, adjusting machine learning to support notions, such as social welfare, that it was not designed to support initially can prove challenging. However, and as we argue throughout, we believe that building on top of existing machinery is a more practical approach than to begin from scratch. The necessity of confronting with welfare consideration can also be an opportunity—as we can leverage these novel challenges to make machine learning practice more informed, transparent, responsible, and socially aware.

### v2

@@ 4.3 Order 2: Enabling User Choices
Even if a system has the capacity to make decisions about users, often those very users will also have some say regarding final outcomes. For example, even in hiring and admissions, candidates must first choose to apply, and if selected, choose to accept. Thus, both reward for the system and utility to users depend on outcomes that result from the interplay of system decisions and user choices. Similarly to system decisions, we will model user choices as also depending on predictions, or more generally on the learned model $h$. Let ${c}_{{\pi}_{h}}(x)$ be the choice of user $x$ in response to $h$. Incorporating user choices into the system’s objective gives:
@@
  • Matching: Prime examples in which user choices drive economic outcomes are recommendation systems and online market platforms. These typically include a single learner whose role is to match users with relevant items (Jagadeesan et al., 2023a); once matched, it is the choices of users that ultimately determine allocations. Since matchings are often based on predictions, their quality determines outcomes for users, suppliers, and the system. This makes them unique in relation to conventional two-sided markets (Rochet and Tirole, 2003). In modern human-facing systems, key resources primarily include the attention, time, and budgets of users. This calls for careful and responsible modeling of user choice behavior (Kleinberg et al., 2024).
  • Competition: For many tasks and services of interest, users can now choose between multiple platforms or providers. This creates competition: once providers aim to maximize their market share, users themselves become the scarce resource, over which providers compete (Ben-Porat and Tennenholtz, 2019; Jagadeesan et al., 2023b; Yao et al., 2023, 2024). When the benefit of users depends on the quality of predictions, competition will revolve around which provider learns more accurate models—and for which users. This creates a novel ‘market for accuracy’, in which learning can have much impact on efficiency and equity.
@@ 5.3 Evaluation
The framework we presented makes use prediction policies (Kleinberg et al., 2015) that specify the dependence of actions on predictions as $a={\pi}_{h}(x)={\pi}(h(x))$. The key assumption of prediction policies is that knowing the true $y$ suffices for making an optimal decision; hence the motivation for replacing $y$ with the prediction ${\hat{y}}=h(x)$. But in realistic settings, $y$ is rarely the only source of uncertainty. Most decision settings have a casual aspect, meaning that actions $a$ can affect outcomes $y$. Hence, there is no ‘true’ $y$ on which we can draw, and the optimal action $a$ (or ‘intervention’) must include an understanding of its causal effects on $y$.
@@ 6 Discussion
Economics studies the tensions that arise from having limited means but unlimited wants. Our main thesis is that while the appetite for what machine learning can give us only grows, we do not yet have the tools to integrate the limitations inherent to any social system in a way that supports informed policy choices. Welfare economics provides general guidelines for designing systems that enable a social planner to express and promote outcomes that align with societal preferences, as they relate to scarcity in resources. We have identified ways in which we believe this perspective applies to, and can be useful for, the design of learning systems that operate in social contexts. Scarcity in learned systems is often far more illusive than in classic economic settings that consider tangible goods. But nonetheless, it exists, and its implications bear much significance for the value that we can hope to obtain from using machine learning as individuals. Our proposal to augment conventional learning frameworks with a notion of social welfare serves as a first step towards this goal. But the road ahead is long and challenging, and requires careful deliberation of the societal role we envision for machine learning.
Acknowledgments
Haifeng Xu is supported by the AI2050 program at Schmidt Sciences (Grant G-24-66104), Army Research Office Award W911NF-23-1-0030, and NSF Award CCF-2303372. Nir Rosenfeld is supported by the Israel Science Foundation grant no. 278/22.
Appendix A Example social welfare functions
Welfare economics considers a wide array of social welfare function types. Here we survey a small subset of the most popular alternatives. The primary distinction is between ordinal functions, which rank all possible economic states in terms of their social preference, and cardinal functions, which assign particular scalar values to states. While any cardinal function induces an ordering, instantiating such functions typically requires access to (and the existence of) individual utility values, which are not always observed (or even attainable) and can be compared on a single shared scale (e.g., monetary value). This is why much of the literature considers cardinal functions, as we do here.
The simplest form of social welfare function considers the average utility over all individuals in a society:
$\frac{1}{n}\sum_{i=1}^{n}u_{i}$
which is known as utilitarian welfare. This is commonly generalized to include (i) a mapping $g$ of individual utilities $u_{i}$ to socially relevant values, and (ii) a weighting scheme $w$ that assigns different weights to different cases. The general form is:
$\frac{1}{n}\sum_{i=1}^{n}w_{i}g(u_{i})$
where $g(u)=u$ and $w_{i}=1$ recovers pure utilitarian welfare. A common choice is setting $g$ to be concave to encodes risk aversion. Weights can more generally be functions (rather than fixed scalar), and both $g$ and $w$ can depend on individual utilities or on the set of all utilities, on individual properties (features) $x_{i}$, and on additional exogenous variables $z$. Thus, despite its simple form, average welfare can be a rather flexible notion. These extensions apply to all social welfare functions described below, and so for simplicity we will present them in the above format.
An alternative to prioritizing is the egalitarian approach which instead aims to maximize welfare for the worst-off individuals in society. This is known as Rawlsian welfare, given by:
$\min\nolimits_{i\in[n]}w_{i}g(u_{i})$
Note this can be expressed also by an average welfare function with utility-dependent weights defined as $w_{i}(u)=\mathds{1}{\{{i=\operatorname*{argmin}_{j}g(u_{j})}\}}$ for $u=(u_{1},\dots,u_{n})$. This suggests a smooth transition between the two types using softmin weights, $w_{i}\propto e^{-\beta g(u_{i})}$, which as a vector approaches the argmin indicator for $\beta\to\infty$. Less common, but possible, is to replace the min with a max to promote the most well-off individual (with no consideration of others); this applies for example to settings in which different societies compete in winner-takes-all matches, and the goal is to advance the best possible candidate (e.g., international sports competitions).
Other common alternatives exist for interpolating between the utilitarian and Rawlsian approaches. One approach is to consider a convex combination of both, namely:
$\alpha\frac{1}{n}\sum_{i=1}^{n}w_{i}g(u_{i})+(1-\alpha)\min_{i\in[n]}w_{i}g(u_{i})$
for some $\alpha\in[0,1]$. A second is to use geometric averaging, known as Nash welfare:
$\left(\prod_{i=1}^{n}w_{i}g(u_{i})\right)^{\frac{1}{n}}$
The third and most flexible approach is the iso-elastic parameterization:
$\frac{1}{n(1-\gamma)}\sum_{i=1}^{n}(w_{i}g(u_{i}))^{1-\gamma}$
which recovers utilitarian welfare for $\gamma=0$, Nash for $\gamma\to 0$, and Rawlsian for $\gamma\to\infty$.
Social welfare functions can also be made to explicitly account for inequality. For example, Sen’s welfare function is $W(u)(1-G(u))$ where $W$ is any base welfare function, and $G$ is the Gini inequality index. Foster’s welfare similarly considers $W(u)e^{-T(u)}$ where $T$ is the Theil index.